### v3

Abstract
Accurate interpretation of numerical data in financial reports is critical for markets and regulators. Although XBRL (eXtensible Business Reporting Language) provides a standard for tagging financial figures, mapping thousands of facts to over 10k US-GAAP concepts remains costly and error-prone. Existing benchmarks oversimplify this task as flat, single-step classification over small subsets of concepts, ignoring the hierarchical semantics of the taxonomy and the structured nature of financial documents. Consequently, these benchmarks fail to evaluate Large Language Models (LLMs) under realistic reporting conditions. To bridge this gap, we introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first comprehensive benchmark for structure-aware and full-scope XBRL tagging. We decompose the complex tagging process into two subtasks: (1) FinNI (Financial Numeric Identification), which extracts entities and types from heterogeneous contexts (text and tables); and (2) FinCL (Financial Concept Linking), which maps extracted entities to the full US-GAAP taxonomy. This two-stage formulation enables a fair assessment of LLMs’ capabilities in numerical reasoning and taxonomy alignment. Evaluating diverse LLMs in zero-shot settings reveals that while models generalize well in extraction, they struggle significantly with fine-grained concept linking, highlighting critical limitations in domain-specific structure-aware reasoning. Code is available at GitHub  [^1] https://github.com/The-FinAI/FinTagging and datasets are available at Hugging Face  [^2] https://huggingface.co/collections/TheFinAI/fintagging.
0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:: Benchmarking LLMs for Extracting and Structuring Financial Information
Yan Wang1, Lingfei Qian1, Xueqing Peng1, Yang Ren1, Yi Han2, Keyi Wang3, Dongji Feng4, Fengran Mo5, Shengyuan Lin6, Qinchuan Zhang7, Kaiwen He3, Chenri Luo3, Jianxing Chen3, Junwei Wu3, Chen Xu1, Ziyang Xu1, Jimin Huang1thanks:  Corresponding author, Guojun Xiong8, Xiao-Yang Liu3, Qianqian Xie1, Jian-Yun Nie5 1The Fin AI, 2Georgia Institute of Technology, 3Columbia University, 4California State University 5University of Montreal, 6Carnegie Mellon University, 7Rensselaer Polytechnic Institute, 8Harvard University Correspondence: jimin.huang@thefin.ai
1 Introduction
Although LLMs excel at question answering qian2025fino1, information extraction balasubramanian2025leveraging, and long-document summarization yuan2025strucsum, their ability to perform financial tagging remains underexplored. Unlike open-ended generation, tagging requires models to commit to a precise semantic interpretation for each reported number, grounding it in dispersed evidence across tables and surrounding text, and aligning it to a standardized concept that downstream systems can interpret consistently. Each year, over 2 million companies publish financial reports disclosing earnings, expenses, and liabilities, providing essential information for investors, regulators, and analysts. However, inconsistent terminology complicates reliable interpretation. XBRL, introduced in 1999, defines the global standard for machine-readable financial reporting richards2006introduction and is now adopted in 65 countries [^3] https://www.xbrl.org/the-standard/what/what-is-xbrl/. In practice, XBRL tagging requires grounding thousands of reported numbers per filing in mixed table and text evidence and aligning them to over 17,000 standardized taxonomy concepts(Figure 1), a workflow that remains largely manual and error-prone [^4] https://xbrl.us/wp-content/uploads/2023/03/DQC-SECMeetingNotes-20240314.pdf.
Figure 1: An example of Financial Tagging in realistic scenarios, where the blue numbers mark facts to be tagged.
Nevertheless, current XBRL tagging benchmarks often emphasize what is easiest to annotate rather than what is hardest to understand. Many reduce tagging to classification over a truncated taxonomy and remove table structure that carries substantial semantic evidence in real filings (Table 1). As a result, existing datasets typically cover only a few thousand common concepts instead of the full 17,000+ taxonomy, which limits evaluation on rare but important concepts and makes it difficult to assess generalization under long-tailed label distributions. In addition, prior work commonly formulates tagging as closed-set extreme classification loukas2022finer; sharma2023financial. This framing places heavy weight on recalling label identifiers and becomes brittle when the full taxonomy cannot be represented in-context, increasing the likelihood of invalid or hallucinated tags. Consistent with this concern, our ablations show near-zero performance even for strong LLMs such as DeepSeek-V3 liu2024deepseek and GPT-4o hurst2024gpt under full-taxonomy settings. Finally, existing benchmarks focus primarily on narrative text and omit tables, despite tables being the main carrier of numerical disclosures. Consequently, they do not evaluate the joint reasoning over multi-level tabular structure and surrounding text that has been studied in financial QA chen2021finqa; zhu2021tat and is required for realistic end-to-end tagging.
Table 1: Detailed comparison of financial NLP benchmarks across task types, sources, and structural capabilities. “Num.Reasoning” indicates the numerical reasoning. “Struct.IE” denotes the structured information extraction.
Benchmark Scenario Data Source Task Modality #Entity Label #Taxonomy Label Num. Reasoning? Struct. IE? Concept Linking?
FinQA chen2021finqa Decision making SEC 10-K QA text/table 0 0
ConvFinQA chen2022convfinqa Decision making SEC 10-K QA text/table 0 0
TAT-QA zhu2021tat Financial analysis Chinese financial reports QA text/table 0 0
DocVQA mathew2021docvqa Enterprise automation Financial document images QA text/image 0 0
FiNER-ORD shah2023finer Financial tagging Chinese financial disclosures Classification text 49 0
FinRED sharma2022finred Knowledge graph construction English financial news Classification text 38 0
FiNER loukas2022finer XBRL tagging SEC 10-K Extreme classification text 0 139 / 17,688
FNXL sharma2023financial XBRL tagging SEC 10-K Extreme classification text 0 2,800 / 17,688
FINTAGGING Financial & XBRL tagging SEC 10-K IE + Alignment text/table 5 17,688 / 17,688
To address these issues, we introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first LLM-oriented benchmark that evaluates whether models can recover the meaning of financial facts by extracting numerical values from mixed table and text inputs and aligning them to the full US-GAAP taxonomy. Unlike prior work loukas2022finer; sharma2023financial that formulates tagging as flat classification over a truncated label set, 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: requires end-to-end reasoning over both narrative disclosures and tables, and scales evaluation to the complete 17k+ concept space. 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: adopts a two-stage task formulation that separates decisions typically conflated in extreme classification. Financial Numerical Identification (FinNI) asks models to identify numerical facts together with their supporting evidence, while Financial Concept Linking (FinCL) maps each fact to the correct US-GAAP concept by semantic comparison with taxonomy definitions, reframing tagging as alignment rather than label recall. We release two expert-verified evaluation sets, FinNI-eval and FinCL-eval, derived from real-world XBRL filings, along with a unified protocol that evaluates both extraction accuracy and concept-alignment performance.
Across 13 state-of-the-art LLMs, our results show a consistent extraction–linking gap. While top models reliably recover numerical facts from mixed table and text inputs, they struggle to assign the correct US-GAAP concepts, indicating that representations sufficient for extraction do not consistently encode the latent semantic conditions that distinguish closely related concepts, such as scope, aggregation level, and qualifying context. As a result, FinNI performance is strong, whereas FinCL remains below 0.19 even after candidate retrieval, with most errors arising from near-neighbor concepts that share similar surface cues but differ in these underlying conditions. Pipeline controls further clarify this failure mode: accuracy improves when gold entities are provided and improves again when the gold concept is included among candidates, yet it does not converge, showing that errors persist even when evidence is localized and the correct concept is available. The restricted cash example illustrates this pattern, where the model selects a narrower concept triggered by a salient local phrase, despite broader contextual evidence supporting an aggregate concept.
We conclude our main contributions as follows: (1) A New Formulation: We introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first benchmark that reframes XBRL tagging into a two-stage “Extract-and-Align” paradigm, enabling full-scope evaluation against 17k+ US-GAAP concepts, a scale previously intractable for zero-shot LLMs. (2) High-Fidelity Resources: We release two expert-verified datasets, FinNI-eval and FinCL-eval, which incorporate complex tabular structures to assess joint reasoning across structured and unstructured financial modalities. (3) Diagnostic Insights: By benchmarking 13 SOTA LLMs, we identify a critical “knowledge-alignment gap”, demonstrating that while LLMs excel at numeric extraction, they require our alignment framework to prevent performance collapse during large-scale taxonomy mapping.
2 Related Work
From XBRL Classification to Alignment. Traditional XBRL tagging has been framed as extreme multi-class classification. Benchmarks such as FiNER loukas2022finer and FNXL sharma2023financial evaluate models on truncated label sets (139–2,800 tags), and methods like GalaXC saini2021galaxc, SECBERT loukas2022finer, and AttentionXML sharma2023financial focus on coping with label sparsity through hierarchical modeling. However, these closed-set formulations overlook the semantic reasoning required to navigate the full 17k+ US-GAAP taxonomy and are poorly matched to LLM capabilities. Recent LLM-based tools such as XBRL-Agent han2024xbrlagent explore filing analysis but do not evaluate end-to-end fact-to-taxonomy alignment. 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: addresses this gap by reframing tagging as a two-stage extract-and-align task.
Financial Document Understanding. In parallel, financial NLP has progressed from entity and relation extraction (e.g., FiNER-ORD shah2023finer, FinRED sharma2022finred) to numerical reasoning over documents. Benchmarks such as FinQA chen2021finqa, TAT-QA zhu2021tat, and DocVQA mathew2021docvqa require joint reasoning over text and tables. Yet these tasks are designed for QA and do not constrain outputs to standardized accounting concepts. Our benchmark instead requires both multi-modal fact extraction and precise alignment to official US-GAAP definitions, focusing on the “last mile” of structured financial reporting.
3 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
3.1 Task Formulation
Formally, given a financial document $D$ consisting of multiple textual contents and structured tables, a set of predefined value data types $L$ specified by the XBRL specification, and a taxonomy database $\mathcal{T}$ containing a set of financial semantic concepts, the XBRL tagging task is to identify all relevant numerical values in $D$ and annotate each with a structured triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$. Here, Tag denotes the linked concept in $\mathcal{T}$ (e.g. “us-gaap:CashAndDueFrom Banks”, etc.), Fact represents the extracted numerical value as it appears in context (e.g., “2.5”, “10,000”, “two”, etc.), and Type specifies the corresponding data type in $L$ (including monetaryItemType, percentItemType, sharesItemType, perShareItemType, and integerItemType). Each numerical fact is assigned a single, most specific tag from the taxonomy, without overlapping. Specifically, monetaryItemType refers to a financial amount, such as revenue or expenses, typically reported in currency units (e.g., USD). percentItemType represents a rate or ratio, expressed as a decimal (e.g., 25%). sharesItemType indicates the number of shares, such as stocks held or issued. perShareItemType captures values reported on a per-share basis, like earnings per share. Lastly, integerItemType is used for whole number counts, such as the number of employees or transactions.
We formalize the task as a mapping:
$f:(D,L,\mathcal{T})\mapsto\left\{(\texttt{Fact}_{i},\texttt{Type}_{i},\texttt{Tag}_{i})\right\}_{i=1}^{n}$ (1)
where each triplet corresponds to a financial value mention extracted from $D$ and semantically grounded in $\mathcal{T}$.
Inspired by information extraction and alignment works wu2019scalable; wadden2019entity; mukherjee2022ectsum; wang2022conditional and after discussing with financial reporting specialists, we formulated 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: into two sub-tasks: Financial Numeric Identification (FinNI), a multi-modal numerical information extraction task, detects numerical value in $D$ and classifies each with its appropriate Type. Financial Concept Linking (FinCL), a numerical entity normalization task, then associates each identified value with its most appropriate Tag in $\mathcal{T}$ based on contextual and structural cues.
Financial Numeric Identification (FinNI)
The first subtask of 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: focuses on identifying numerical values in a financial document and assigning each a coarse-grained value data type. This corresponds to detecting the Fact and Type components of each triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$ defined in the overall task.
We formalize this subtask as a mapping:
$f_{\text{FinNI}}:(D=(S,T),\ L)\mapsto\left\{(e_{i},l_{i})\right\}_{i=1}^{k}$ (2)
where $S$ and $T$ represent the textual and tabular components of the document, respectively, and $L$ is the set of predefined value data types. Each $e_{i}$ is a numerical entity extracted from either $S$ or $T$, and $l_{i}\in L$ denotes its assigned data type.
Financial Concept Linking (FinCL)
Building on the output of the FinNI subtask, the goal of FinCL is to semantically ground each identified numerical entity $e$ by linking it to a concept $\hat{c}$ in a predefined financial taxonomy, which is the Tag component of each triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$ defined in the overall task. Formally, we define the mapping as:
$f_{\text{FinCL}}:(e,l,C_{e},\mathcal{T})\mapsto\hat{c}$ (3)
where $e$ is a numerical entity extracted from $D=(S,T)$, and $l\in L$ is its associated data type provided by the FinNI stage. $C_{e}$ represents the structural context of $e$, which encompasses linguistic dependencies for narrative mentions and hierarchical row-column coordinates for tabular mentions. $\mathcal{T}=\{c_{1},c_{2},\dots,c_{n}\}$ denotes the full-scope US-GAAP taxonomy containing $n$ uniquely defined and semantically grounded concepts. The objective of FinCL is to identify the optimal concept $\hat{c}\in\mathcal{T}$ that maximizes the semantic alignment between the fact $e$ and its specific context $C_{e}$ within the high-dimensional taxonomy space.
3.2 Raw Data Collection
We compiled 142 annual 10-K filings from publicly listed U.S. companies submitted in 2023–2024 (Table 2). To ensure coverage diversity, the collection spans all 11 major industry sectors, with about 95% of companies located across 31 U.S. states and the remainder based internationally. Using BeautifulSoup, we parsed the reports and extracted 319,893 narrative sentences and 21,576 financial tables. Company-level statistics are provided in Appendix C.1. All filings follow the SEC Inline XBRL (iXBRL) format, which embeds machine-readable tags inside HTML disclosures and links textual and tabular values to US-GAAP taxonomy concepts through schema and linkbase definitions. However, these regulatory tags are designed for compliance rather than model evaluation. They provide little coverage of negative cases and contain substantial structural noise. We therefore construct a high-fidelity evaluation suite that repurposes this raw regulatory metadata into a standardized benchmark for financial fact extraction and concept linking. Additional discussion of our benchmark granularity is provided in Appendix C.2.
Table 2: Statistics of collected financial reports.
Item Value
Report type 10-K
Period 2023-02-13 to 2025-02-13
#Company 142
#Covered sector 11
#Covered jurisdiction 31 States + outside US
#Sentence 319,893
#Char 75,748,949
#Table 21,576
3.3 Large-Scale Automated Annotation
Our automated pipeline produces a large, high-fidelity corpus reflecting real financial disclosures. As summarized in Table 3, the dataset contains 15,986 narrative sentences and 12,801 tables, yielding 261,457 numerical entities linked to 3,953 unique US-GAAP concepts. Including both valid facts and naturally occurring noise, it supports a realistic evaluation of LLMs under regulatory-style conditions.
Table 3: Statistical information for the original dataset in our benchmark. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
Item Sentence Table
Positive instances 7,768 8,709
Negative instances 8,218 4,092
Avg. Tokens/S $80.91\pm 63.62$ $1212.42\pm 1421.76$
Avg. Entities/S $1.24\pm 1.82$ $18.87\pm 37.16$
Avg. Concepts/S $1.24\pm 1.82$ $18.87\pm 37.16$
Total Entities 261,457
Entity Types 5
Unique Concepts 3,953
Automated Annotation Engine. We build a rule-based parser (Algorithm 1, Appendix C.3) that reconstructs semantic links between human-readable disclosures and Inline XBRL (iXBRL) metadata. For each instance $i$, the parser locates ix:nonfraction elements and extracts the value $v$, concept $c$, and data type $\tau$, while restricting extraction to atomic leaf elements $\epsilon$ to avoid duplicates from nested tags.
Valid Entity Types. To define the label space, we profile all numeric XBRL item types. As shown in Figure 2, the distribution is long-tailed: although eleven types appear, five dominate. We therefore restrict $\mathcal{L}$ to monetaryItemType, percentItemType, sharesItemType, perShareItemType, and integerItemType. An instance is labeled POSITIVE ($\mathcal{P}$) only if it contains at least one concept $c\in\mathcal{T}$ from $\mathcal{L}$; otherwise it is labeled NEGATIVE ($\mathcal{N}$).
Structural Refinement and Filtering. For reliable evaluation, we keep only text segments longer than 20 characters and deduplicate content globally using $\mathcal{S}_{text}$. For tables, we retain structural tags (<table>, <tr>, <td>, <th>) but discard tables without numerical facts. This filtering preserves semantic accuracy while maintaining realistic layout structure.
Figure 2: Statistics of numerical entity types.
3.4 Data Quality Control
Validating all 261,457 entities is infeasible. To obtain a rigorous quality upper bound, we adopt a strategic sampling scheme based on a greedy minimal-cover algorithm (Algorithm 2, Appendix D.1). Rather than random sampling, we construct a diversity-maximized subset $\mathcal{P}^{*}$ that satisfies two goals: (1) Company Balance, limiting bias from individual reporting styles, and (2) Taxonomy Completeness, ensuring that each US-GAAP concept observed in the corpus appears at least once. We seed $\mathcal{P}^{*}$ with capped samples per company, then iteratively add instances that cover the largest number of uncovered concepts, producing a compact but representative audit set. The final subset contains 928 positive sentence instances and 1,021 positive table instances, serving as a taxonomy-wide stress test reviewed by two junior analysts and one senior adjudicator (the demography of these annotators is shown in Appendix D.2).
Review Protocol. The audit followed two stages. A pilot phase aligned interpretations of the Verification Guideline (Appendix D.3), after which analysts independently verified whether extracted triplets $(e,\tau,c)$ were fully correct. The senior expert adjudicated disagreements to create the final gold set.
Agreement Results. On the independent round, annotators achieved 96% raw agreement and Cohen’s $\kappa=0.81$, indicating substantial agreement landis1977measurement, the equation is shown in Appendix D.4. This high-verified accuracy on the most semantically diverse subset provides strong evidence for the reliability of our automated annotation pipeline.
3.5 Task-Specific Dataset Construction
Following validation, we integrated the corrected $\mathcal{P}^{*}$ subset back into the full corpus to form a high-fidelity hybrid dataset. This refined pool ensures that every US-GAAP concept in the taxonomy has at least one expert-verified representation. From this foundation, we derive two subtask-specific datasets: FinNI-eval, targeting numerical entity identification, and FinCL-eval, targeting concept linking. Both are the first datasets in the financial domain to emphasize structured information extraction and concept-level alignment. Their detailed statistics are reported in Table 4.
Table 4: Statistics of evaluation datasets for FinNI-eval and FinCL-eval. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
Item FinNI-eval FinCL-eval
#Instance 28,787 261,457
Avg. Input Tokens $1101.50\pm 992.00$ $60.43\pm 48.25$
Max Input Tokens 18,831 750
Avg. Output Tokens $175.24\pm 375.90$ $12.82\pm 5.09$
Max Output Tokens 7,813 41
3.5.1 FinNI-eval Dataset Construction
Based on the annotation procedure described in Section 3.3, we construct the FinNI-eval dataset, comprising 28,787 instances derived from the annotated sentences and tables. Specifically, the dataset comprises two components: the input block and the answer block. The input block includes a task instruction and the input content. The instruction defines the FinNI task by specifying the responsibilities of the LLMs, the definitions of entity types, and the rules to be followed. The answer block is formatted as a list aligned with the input block. The answer is a JSON list of all identified entity values and types or an empty JSON list. The specific prompt template is shown in Figure 9.
FinNI-eval Dataset
Instruction: <FinNI Task Instruction>
Input: <Sentences or Table>
Answer: {"results": [{"Fact": <numeric entity>,
        "Type": <entity type>}]} or {"results": []}
3.5.2 FinCL-eval Dataset Construction
Following the formulation of the FinCL subtask in Section 3.1, we further construct a FinCL-eval dataset that includes 261,457 query-answer pairs for numerical entity normalization. In addition, we built a US-GAAP taxonomy database containing 17,688 unique financial concepts. To create the FinCL-eval dataset, we utilize all positive instances identified in Section 3.3. Each query consists of a numerical entity, its corresponding entity type, and its surrounding context, while the answer is the associated US-GAAP concept.
FinCL-eval Dataset
Query: <entity> + <entity type> + <context>
Answer: <US-GAAP Concept>
3.6 Evaluation
3.6.1 Evaluation Framework
We propose a unified evaluation framework for 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:. As illustrated in Figure 3, each input is decomposed into two subtasks, FinNI and FinCL, which together produce structured triplets $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$. This design jointly evaluates zero-shot fact extraction and concept alignment. For comparison with token-classification baselines, we report macro and micro Precision, Recall, and F1 sokolova2009systematic. FinNI is evaluated using pair-level metrics, and FinCL with accuracy. Full metric definitions are in Appendix E.
Figure 3: A unified evaluation framework on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark. Note: The Input content is drawn from the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: dataset, and the “Shared LLMs” are used for identification in the FinNI task and reranking in the FinCL task.
FinNI Evaluation. We adopt pair-level metrics to measure extraction quality, focusing on whether models can locate financial values from mixed text–table contexts based on semantic cues.
FinCL Evaluation. To test fine-grained numeric entity normalization, we formulate FinCL as a retrieval–reranking task: a retrieval stage produces a candidate set of taxonomy concepts, and the LLM then reranks candidates to select the most appropriate tag, avoiding impractical extreme classification (prompt template in Figure 10, Appendix).
To this end, we embed each taxonomy concept using text-embedding-3-small [^5] https://platform.openai.com/docs/models/text-embedding-3-small , then retrieve the top-$h$ candidate tags from the US-GAAP taxonomy $\mathcal{T}$ by semantic similarity. The LLM subsequently reranks these candidates using contextual reasoning to select the best tag $\hat{c}\in\mathcal{T}_{h}$:
$f_{\text{rerank}}:(e,l,C_{e},\mathcal{T}_{h})\mapsto\hat{c}$ (4)
3.6.2 Evaluation Models
Our goal is to assess the foundational capabilities of SOTA LLMs on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and understand their strengths and limitations in XBRL tagging. We evaluate models across three categories using our unified framework (Table 11). (1) Closed-source general LLM: GPT-4o hurst2024gpt. (2) Open-source general LLMs: DeepSeek-V3 liu2024deepseek, DeepSeek-R1-Distill-Qwen-32B deepseekai2025deepseekr1incentivizingreasoningcapability, Qwen3-32B / 14B / 1.7B / 0.6B qwen3technicalreport, Llama-4-Scout-17B-16E-Instruct meta2025llama, Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct grattafiori2024llama, and gemma-2-27b-it gemma_2024. (3) Domain-specific financial LLM: Fino1-8B qian2025fino1. We also compare against strong PLM baselines: BERT-large DBLP:journals/corr/abs-1810-04805, FinBERT araci2019finbert, and SECBERT loukas2022finer.
3.6.3 Evaluation setting
[…]
4 Experiment and Result
4.1 Overall Results
Table 5 presents the overall performance on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark. It clearly demonstrates that under our framework, LLMs can effectively handle both frequent and rare financial tags, indicating their ability to mitigate long-tail label challenges and underscoring the advantage of our information extraction and alignment formulation over traditional token-level classification approaches.
Table 5: Overall Performance. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
Category Models Macro Micro
P R F1 P R F1
Closed-source LLM GPT-4o 0.0865 0.0626 0.0569 0.1052 0.0731 0.0863
Open-source LLMs DeepSeek-V3 0.0949 0.0778 0.0659 0.1074 0.1264 0.1162
Llama-4-Scout-17B-16E-Instruct 0.0683 0.0414 0.0400 0.1045 0.0526 0.0699
Llama-3.3-70B-Instruct 0.0544 0.0279 0.0288 0.0665 0.0382 0.0485
DeepSeek-R1-Distill-Qwen-32B 0.0532 0.0283 0.0285 0.0814 0.0214 0.0339
Qwen3-32B 0.0639 0.0314 0.0324 0.1127 0.0230 0.0382
gemma-2-27b-it 0.0471 0.0291 0.0276 0.0533 0.0390 0.0451
Qwen3-14B 0.0591 0.0274 0.0288 0.1069 0.0182 0.0311
Llama-3.1-8B-Instruct 0.0345 0.0178 0.0169 0.0575 0.0166 0.0258
Llama-3.2-3B-Instruct 0.0194 0.0118 0.0100 0.0178 0.0084 0.0114
Qwen3-1.7B 0.0207 0.0067 0.0080 0.1063 0.0031 0.0060
Qwen3-0.6B 0.0026 0.0008 0.0010 0.0562 0.0002 0.0005
Financial LLM Fino1-8B 0.0344 0.0143 0.0151 0.0419 0.0128 0.0197
Fine-tuned PLMs BERT-large 0.0252 0.0266 0.0205 0.1518 0.1283 0.1391
FinBERT 0.0046 0.0064 0.0042 0.0872 0.0526 0.0656
SECBERT 0.0231 0.0295 0.0203 0.1870 0.1697 0.1779
From a macro perspective, which emphasizes balanced performance across frequent and rare tags, large general-purpose LLMs clearly dominate. DeepSeek-V3, GPT-4o, and Llama-4-Scout achieve the strongest macro-F1 scores, surpassing all fine-tuned PLMs and indicating better generalization to long-tail concepts. Qwen3-32B also performs competitively, suggesting that model architecture and pretraining can compensate for a smaller scale. From a micro perspective, which reflects performance on frequent labels, the trend is similar: top LLMs remain competitive even without fine-tuning, while fine-tuned PLMs such as SECBERT and BERT-large excel mainly because they benefit from label frequency rather than broader coverage (Table 5).
4.2 Subtask Results
4.2.1 FinNI subtask
Table 6 summarizes results on FinNI. Overall, larger LLMs show markedly stronger ability to identify numerical entities and produce structured outputs, even without financial adaptation. DeepSeek-V3 achieves the best overall performance, while GPT-4o performs competitively and ranks among the top models.
Table 6: Performance comparison of different models on the FinNI subtask. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
Category Model Precision Recall F1
Closed-source LLM GPT-4o 0.5893 0.4977 0.5397
Open-source LLMs DeepSeek-V3 0.5886 0.8430 0.6932
Llama-4-Scout-17B-16E-Instruct 0.4668 0.3164 0.3771
Llama-3.3-70B-Instruct 0.4826 0.3301 0.3920
DeepSeek-R1-Distill-Qwen-32B 0.5676 0.1942 0.2894
Qwen3-32B 0.6991 0.1804 0.2868
gemma-2-27b-it 0.5060 0.4526 0.4778
Qwen3-14B 0.6912 0.1487 0.2448
Llama-3.1-8B-Instruct 0.3874 0.1761 0.2421
Llama-3.2-3B-Instruct 0.1856 0.1203 0.1460
Qwen3-1.7B 0.7362 0.0281 0.0541
Qwen3-0.6B 0.2803 0.0019 0.0038
Financial LLM Fino1-8B 0.3431 0.1293 0.1878
A notable pattern is that some models (e.g., Qwen3 variants) obtain very high precision but extremely low recall, indicating that they extract only a small fraction of target facts. This confirms that precision alone is insufficient for FinNI, where missing facts are often more damaging than producing a few spurious ones.
Smaller open-source models struggle severely, reflecting the difficulty of this zero-shot extraction setting. In contrast, mid-size models such as gemma-2-27B achieve more balanced behavior. Interestingly, the financial LLM Fino1-8B still lags behind general-purpose models, suggesting that domain text pretraining alone does not help unless the training task aligns closely with FinNI.
4.2.2 FinCL subtask
Table 7 reports FinCL performance under the reranking setup, using candidates retrieved from the best top-200 configuration (Table 14). Although retrieval provides essential coverage, low Acc@k values, especially for table entities, substantially cap downstream accuracy.
Table 7: Performance comparison of different models on the FinCL subtask. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
Category Model Accuracy
Closed-source LLM GPT-4o 0.1829
Open-source LLMs DeepSeek-V3 0.1889
Llama-4-Scout-17B-16E-Instruct 0.1649
Llama-3.3-70B-Instruct 0.1318
DeepSeek-R1-Distill-Qwen-32B 0.1141
Qwen3-32B 0.1277
gemma-2-27b-it 0.1099
Qwen3-14B 0.1144
Llama-3.1-8B-Instruct 0.0913
Llama-3.2-3B-Instruct 0.0415
Qwen3-1.7B 0.0735
Qwen3-0.6B 0.0414
Financial LLM Fino1-8B 0.0807
Within this setting, large LLMs such as DeepSeek-V3, GPT-4o, and Llama-4-Scout perform best, but overall accuracy remains low, reflecting the difficulty of aligning evidence with fine-grained taxonomy concepts. Most other models cluster well below this tier, and smaller models approach random behavior. The financial model Fino1-8B also lags behind general LLMs, suggesting that domain pretraining alone is insufficient without task-aligned supervision.
Taken together, FinCL exposes two compounding challenges: taxonomy disambiguation requires reasoning over subtle semantic cues, and retrieval errors propagate directly into reranking, creating a ceiling effect on end-to-end alignment.
4.3 Ablation analysis
We further compare our benchmark against extreme multi-class classification settings. The prompt template is shown in Figure 11 in the Appendix, and we select the best-performing model from each category. The evaluation uses triplet-level ($\{\texttt{Tag},\texttt{Fact},\texttt{Type}\}$) Precision, Recall, and F1.
Table 8: Performance comparison between with/without our evaluation framework on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark dataset.
Evaluation Mode Model Precision Recall F1
FINTAGGING GPT-4o 0.1166 0.0915 0.1026
DeepSeek-V3 0.1187 0.1581 0.1356
Fino1-8B 0.0336 0.0140 0.0197
Extreme Classification GPT-4o 0 0 0
DeepSeek-V3 0 0 0
Fino1-8B 0 0 0
As shown in Table 8, treating XBRL tagging as single-step extreme classification fails in practice. The model must choose directly from thousands of taxonomy concepts without first identifying valid numerical entities, forcing it to rely on internal priors rather than grounded evidence. All models collapse to zero, illustrating that this formulation is not a meaningful way to evaluate LLMs on XBRL tagging. In contrast, our two-stage framework mirrors how tagging is actually performed. FinNI first detects and types candidate numerical facts, and FinCL then retrieves a small set of plausible tags and requires the model to select among them using context. Under this formulation, models achieve non-trivial performance, revealing real differences in capability that are completely hidden under extreme classification.
4.4 Error propagation in 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: pipeline
Table 9 compares accuracy across three settings: the full pipeline, FinCL with gold entities, and rerank-only with guaranteed gold candidates. The pattern reveals clear cascading errors. End-to-end performance is consistently the lowest because mistakes from extraction and retrieval accumulate in the linking stage.
When gold entities are supplied, accuracy improves substantially, showing that entity identification is a major bottleneck. Performance increases again in the rerank-only setting, indicating that narrowing the label space reduces difficulty. However, accuracy still remains far from solved, suggesting that fine-grained US-GAAP disambiguation is intrinsically challenging even when retrieval is perfect. In addition, we provide representative error cases in Appendix I, illustrating how these challenges arise in practice.
Table 9: Accuracy comparison across the full pipeline, FinCL subtask, and rerank-only settings.
Model Pipeline FinCL Rerank-only
GPT-4o 0.1052 0.1829 0.2023
DeepSeek-V3 0.1074 0.1889 0.2144
Llama-4-Scout-17B-16E-Instruct 0.1045 0.1649 0.1792
Llama-3.3-70B-Instruct 0.0665 0.1318 0.1482
DeepSeek-R1-Distill-Qwen-32B 0.0814 0.1141 0.1258
Qwen3-32B 0.1127 0.1277 0.1402
gemma-2-27b-it 0.0533 0.1099 0.1210
Qwen3-14B 0.1069 0.1144 0.1321
Llama-3.1-8B-Instruct 0.0575 0.0913 0.1076
Llama-3.2-3B-Instruct 0.0178 0.0415 0.0566
Qwen3-1.7B 0.0598 0.0735 0.0880
Qwen3-0.6B 0.0362 0.0414 0.0533
Fino1-8B 0.0419 0.0807 0.0908
5 Conclusion
This paper introduces 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a benchmark for evaluating LLMs on XBRL tagging of real financial reports. The task is structured into two subtasks, Financial Numerical Identification and Financial Concept Linking, to separately assess extraction and concept alignment. Our results show that LLMs perform reasonably well in zero-shot settings and generalize to long-tail entities, but still struggle to align facts with precise US-GAAP concepts. These findings point to gaps in structure-aware and semantic reasoning. 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: provides a foundation for future work on reliable XBRL tagging and regulatory reporting.
Limitations
In this work, we propose 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and conduct a comprehensive analysis of LLM performance on the task. However, several limitations remain. First, we have not yet evaluated the latest models, such as GPT-5.2, gpt-oss, and Qwen3-235B. We are currently conducting these experiments and will report the results in future work. Second, we collected only a small dataset for fine-tuning PLMs and did not construct task-specific training data for LLMs. Future work may explore building LLM-based XBRL tagging agents through targeted fine-tuning to enhance performance further.
Potential Risks
Potential Positive Societal Impacts.
0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: provides the first structure-aware benchmark for evaluating large language models on realistic XBRL tagging. By reframing tagging as a transparent, two-stage Extract–Align process, this benchmark:
  • Enables more rigorous and reproducible evaluation of model performance on financial fact extraction and taxonomy alignment.
  • Supports research on safer automation of financial workflows by exposing failure modes that arise in long-tailed concept spaces and complex tabular contexts.
  • Lowers the barrier for auditing, interpreting, and improving financial AI systems by releasing open datasets, evaluation protocols, and diagnostic analyses.
By emphasizing open evaluation rather than proprietary pipelines, 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: contributes to more transparent and accountable financial AI research.
Potential Negative Societal Impacts.
Despite these benefits, several risks remain:
  • Misuse and Overreliance. High benchmark performance may be mistakenly interpreted as readiness for real-world XBRL tagging, regulatory reporting, or financial decision-making, even though the benchmark does not replace human review or auditing.
  • Bias and Coverage Limitations. The dataset is drawn from U.S. SEC filings and US-GAAP concepts. As a result, reporting practices, sector distributions, and taxonomy frequency patterns may introduce biases and may not generalize to other jurisdictions or accounting standards.
  • Benchmark Narrowness. 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: evaluates extraction and alignment only. Downstream tasks such as aggregation, reconciliation, audit validation, and materiality assessment remain outside the benchmark and may still fail in deployed systems.
Ongoing iteration, external feedback, and careful documentation are necessary to mitigate these concerns.
Ethical Considerations and Usage Disclaimer
The authors assume responsibility for the design and dissemination of 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:. All data used in the benchmark are derived from publicly available SEC filings and contain no intentionally included personal or sensitive information. Dataset construction followed community ethical guidelines for research involving public financial disclosures.
Unless otherwise specified, the datasets and accompanying evaluation resources are released for research and educational purposes. They should not be used as a substitute for professional accounting judgment, financial reporting systems, or regulatory compliance procedures.
The benchmark is designed to evaluate model behavior under controlled conditions. Performance on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: does not imply suitability for production deployment, financial auditing, or automated investment or compliance tasks. Any real-world use must include human oversight, documentation of limitations, and appropriate legal and regulatory review.
We used AI-assisted writing tools to help refine the wording and improve the readability of the manuscript. All scientific claims, dataset design choices, and experimental analyses were developed and verified by the authors.
While reasonable care has been taken to ensure accuracy, the authors provide no warranty regarding completeness or suitability for any particular purpose and disclaim liability for any damages arising from use of the benchmark or associated materials.
Licensing and Terms of Use
We release all benchmark data, annotations, and accompanying resources under the CC-BY 4.0 license, which permits redistribution and adaptation with attribution. All source documents originate from publicly available SEC filings, and we strictly comply with the original terms governing their public distribution. The US-GAAP taxonomy is used in accordance with its publicly documented terms. Our released benchmark does not contain proprietary material beyond what is already publicly accessible, and any derivative use must continue to acknowledge both the original providers and our benchmark.
Appendix A Significance analysis
To further examine the comparative performance across models, we conducted a pairwise significance analysis, with the results summarized in Figure 4. The upper-triangular matrix in SubFigure 4(a) presents the pairwise significance outcomes, where a value of 1 indicates that the row model’s performance is significantly different from the column model. In contrast, 0 indicates that the difference is not statistically significant. The diagonal elements are set to zero, since self-comparisons are not meaningful.
(a) Model-vs-Model significance matrix.
(b) Per-model count of significant pairwise differences.
Figure 4: Visualization of significance analysis across models. (a) shows the pairwise-bootstrap significance matrix, while (b) summarizes how many significant differences each model has relative to others.
As shown in SubFigure 4(b), the number of significant pairwise differences varies across models, reflecting heterogeneous performance behaviors on the evaluated benchmark. The models DeepSeek-R1-Distill-Qwen-32B and Qwen3-32B exhibit the largest numbers of significant differences, followed by GPT-4o, Llama-4, and Qwen3-14B, suggesting that their performance diverges more noticeably from other models. These results may indicate that models within the Qwen family display stronger variability across comparisons, potentially due to differences in training objectives or architectural configurations. In contrast, DeepSeek-V3, Llama-3.3-70B-Instruct, gemma-2-27b-it, and Fino1-8B demonstrate relatively fewer significant differences, implying that their performance remains more stable and comparable to other high-performing models. Overall, these findings suggest that larger or instruction-tuned models tend to produce more consistent outcomes, while models with smaller sizes or domain-oriented tuning show greater variation and clearer statistical separations under significance testing.
Overall, this analysis highlights not only which models outperform or underperform others but also which models behave similarly when subjected to statistical evaluation. The pairwise significance matrix, therefore, provides an interpretable view of model robustness and performance distinctiveness beyond average accuracy metrics.
Appendix B Literature Review
The XBRL provides a comprehensive taxonomy for financial reporting, encompassing thousands of detailed tags corresponding to concepts within financial statements. Applying NER to assign XBRL tags is an emerging yet challenging area.
B.1 XBRL Tagging benchmark
FiNER systematically benchmarked several neural architectures on the finer-139 dataset to address numeric-heavy XBRL tagging loukas2022finer. The initial experiments showed that standard BERT underperforms due to subword fragmentation, then the authors introduced pseudo-token strategies replacing numerals with [NUM] or [SHAPE] tokens to stabilize label assignment across fragmented numeric spans. These strategies, combined with domain-specific pretraining on SEC-BERT, significantly improved tagging performance, reaching 82.1 micro-F1 without the need for computationally expensive CRF layers. Their experiments demonstrated that subword-aware models with numeric-aware pseudo-tokens outperform both word-level BiLSTMs and vanilla BERT, particularly in numeric-heavy contexts, and avoid nonsensical label sequences. FNXL extended this benchmarking paradigm to a much larger label space of 2,794 US-GAAP tags, reframing the task as an extreme classification problem sharma2023financial. They compared the FiNER sequence labeling approach with a two-step pipeline that first identifies numeric spans and then assigns labels using AttentionXML. While FiNER achieved stronger micro-F1 (75.84), reflecting better performance on frequent tags, AttentionXML outperformed FiNER in macro-F1 (47.54), highlighting its strength in predicting infrequent, tail-end labels. FNXL further evaluated both models under a Hits@k setting, confirming that label recommendations from the AttentionXML pipeline could substantially reduce manual effort and maintain high inter-annotator agreement. Together, these benchmarks reveal the need for context-aware reasoning and label-ranking mechanisms in realistic XBRL tagging scenarios.
B.2 XBRL Tagging Methods
The previous studies also explored approaches to address scalability, semantic ambiguity, and reasoning gaps in XBRL tagging to improve performance. Saini et al. saini2021galaxc proposed GalaXC is a graph-based extreme classification framework that jointly learns over document-label graphs with per-label attention across multi-hop neighborhoods. By integrating label metadata and transitive label correlations, GalaXC outperformed leading deep classifiers by up to 18% in micro-F1 on standard benchmarks and achieved 25% gains in warm-start scenarios where partial labels are available. Moreover, Wang et al. wang2023standardizing addressed the practical challenge of custom tag standardization through a semantic similarity pipeline that leverages TF-IDF, Word2Vec, and FinBERT embeddings. Although unsupervised, the method was tested across nearly 200,000 custom tags from SEC filings between 2009 and 2022, and showed strong alignment performance, with vector-based mappings identifying viable standard tag candidates for a substantial proportion of non-compliant elements, offering a low-cost and interpretable solution for downstream financial analysis. Shifting focus from classification to comprehension, XBRL-Agent evaluated the capabilities of large language models to reason over full XBRL reports han2024xbrlagent. The authors introduced two task types, domain taxonomy understanding and numeric reasoning, and found that base LLMs often hallucinated or misinterpreted financial content. To overcome these issues, XBRL-Agent incorporated retrieval-augmented generation (RAG) and symbolic calculators within an LLM-agent framework. The enhanced system achieved a 17% accuracy gain on domain query tasks and a 42% boost on numeric reasoning queries compared to base LLMs, validating the utility of modular tool augmentation. These improvements enabled reliable multi-step reasoning over complex disclosures such as debt instruments and derivative gains, which are difficult to capture using span-level classifiers. Collectively, these works broaden the methodological landscape of XBRL tagging from graph-based label propagation and embedding-based normalization to LLM-driven report analysis and point to a hybrid future where structured priors and reasoning tools jointly support accurate, scalable financial information extraction.
B.3 Financial Evaluation Benchmarks
In parallel to XBRL-specific advances, the financial NLP community has developed comprehensive benchmarks to assess broader capabilities in information extraction, numerical reasoning, and document understanding. FiNER-ORD shah2023finer introduced a high-quality, domain-specific NER dataset annotated over financial news, emphasizing general entity types like persons, organizations, and locations. While not numerically focused like FiNER-139, it highlights the lexical diversity of financial discourse and establishes a strong baseline for testing pretrained and zero-shot LLMs in real-world financial NER scenarios. FinQA chen2021finqa pushed toward explainable QA by pairing expert-written questions with annotated multi-step reasoning programs derived from earnings reports. ConvFinQA chen2022convfinqa extended this challenge to conversational contexts, simulating real-world question flows over sequential financial queries. TAT-QA zhu2021tat focused on hybrid tabular-text reasoning and required models to align cell values and document narratives, often involving aggregation, comparison, and unit-scale interpretation. Pixiu xie2023pixiu introduced a broader evaluation framework by releasing FinMA, a financial LLM instruction-tuned across five tasks, and assessing it on a new benchmark covering sentiment classification, QA, summarization, NER, and stock prediction. BizBench koncel2023bizbench framed financial QA as program synthesis over realistic, multi-modal contexts, integrating reasoning, code generation, and domain knowledge into a single evaluation pyramid. While these benchmarks highlight the growing ability of models to integrate structured and unstructured financial data, they overlook taxonomy-driven fact alignment and do not support the structured output formats required for XBRL tagging.
Appendix C Data and Ticker Information
C.1 The Statistics of the Tickers
Figure 5, Figure 6, and Figure 7 illustrate the distribution of 142 tickers across industry sectors, market capitalization categories, and geographic regions, highlighting the diversity of our collected financial reports. Considering that practical XBRL tagging practices often vary by industry, company size, and legal jurisdiction, we curated a diverse set of companies covering all 11 major industry sectors. The sample maintains a balanced distribution across firm sizes, following market capitalization categories: micro-cap (<\$300M), small-cap (\$300M–\$2B), mid-cap (\$2B–\$10B), large-cap (\$10B–\$200B), and mega-cap (>\$200B). In addition, we incorporated firms from over 30 states and international jurisdictions to capture regional differences in reporting and tagging conventions. This diversity ensures that the benchmark reflects realistic heterogeneity observed in financial disclosures across sectors, scales, and regulatory environments.
Figure 5: Distribution of Tickers by Industry Sector.
Figure 6: Distribution of Tickers by Market Cap Size.
Figure 7: Distribution of Tickers by Geography.
C.2 Data Granularity Justification
Each 10-K report follows the SEC’s XBRL filing standard, where all narrative and tabular disclosures are linked through machine-readable tags defined in the US-GAAP taxonomy. While the XBRL framework provides document-level structure via instance, schema, and linkbase files, the tagging of financial values and concepts occurs at a much finer granularity—typically within individual sentences or table rows that contain numeric facts and their contextual descriptions.
For this reason, our benchmark adopts the sentence and table sequence as the fundamental input unit rather than entire reports. This design choice is motivated by three considerations:
  • Locality of Semantic Scope: Financial facts and their US-GAAP tags are usually expressed within a single sentence or a bounded table region. Modeling at this level aligns the input scope with the tagging scope, reducing noise from unrelated content within long filings.
  • Computational Efficiency: Full 10-K filings often exceed hundreds of pages and contain heterogeneous sections (e.g., risk factors, management discussion, and financial statements). Processing the entire report as one input sequence would exceed model context limits and obscure local numerical relations. Segment-level inputs enable fine-grained supervision and efficient evaluation.
  • Faithful Tag Alignment: Because each numerical entity is explicitly linked to its taxonomy tag in the XBRL instance document, using sentence or table segments ensures a one-to-one correspondence between input content and ground-truth annotations. This granularity preserves the original tagging fidelity of the filings.
Overall, this segmentation strategy allows our benchmark to remain consistent with how XBRL tagging operates in practice (on localized factual statements), while maintaining scalability and interpretability across large collections of filings.
C.3 Automated Data Annotation
Algorithm 1 converts Inline XBRL (iXBRL) filings into structured instances for our benchmark. The procedure operates over both narrative sentences and tables, and assigns each instance a positive or negative label based on whether it contains at least one validated financial entity.
We begin by iterating through all disclosure instances in the corpus and extracting their plain-text content. Very short fragments (length $\leq 20$) and duplicated passages are discarded using a global text set, ensuring that the benchmark contains unique and meaningful contexts.
For each remaining instance, the algorithm traverses all atomic iXBRL elements. From each element, we retrieve the numerical value $v$, its associated concept $c$ from the US-GAAP taxonomy, and the declared data type $\tau$. We retain only elements whose concept is in $\mathcal{T}$ and whose type belongs to the predefined valid numeric set $\mathcal{L}$. Restricting extraction to leaf-level elements avoids double counting values that appear in nested XML structures. A local deduplication set further prevents repeated annotation of identical $(v,c)$ pairs within the same instance.
If at least one valid entity is identified, the instance is recorded as a positive example together with its entity list; otherwise, it is stored as a negative example. Negative instances are intentionally preserved, as they reflect realistic disclosure text that models must learn to ignore. Overall, this process yields a corpus that is faithful to regulatory metadata while remaining structurally clean and suitable for downstream evaluation.
Input: Raw 10-K reports (iXBRL) $\mathcal{D}$; US-GAAP taxonomy $\mathcal{T}$; Set of valid data types $\mathcal{L}$
Output: Positive instances $\mathcal{P}$; Negative instances $\mathcal{N}$
Initialize: $\mathcal{P}\leftarrow\emptyset$, $\mathcal{N}\leftarrow\emptyset$, Global unique text set $\mathcal{S}_{text}\leftarrow\emptyset$;
foreach instance (sentence or table) $i\in\mathcal{D}$ do
 $text\leftarrow$ Extract plain text content from instance $i$;
 if length of $text\leq 20$ or $text\in\mathcal{S}_{text}$ then
    skip to the next instance;
 $E\leftarrow\emptyset$ // List of validated entities in $i$;
 $S_{fact}\leftarrow\emptyset$ // Local fact deduplication set;
 foreach atomic iXBRL element $\epsilon$ within instance $i$ do
      Retrieve value $v$, financial concept $c$, and data type $\tau$ from $\epsilon$;
    if $c\in\mathcal{T}$ and $\tau\in\mathcal{L}$ and $v$ is not empty then
       if $(v,c)\notin S_{fact}$ then
            Add $(v,\tau,c)$ to the entity list $E$;
            Add $(v,c)$ to the local deduplication set $S_{fact}$;
 if $E$ is not empty then
      Add $(text,E)$ to the positive set $\mathcal{P}$;
 else
      Add $(text,\emptyset)$ to the negative set $\mathcal{N}$;
  Add texttext to the global set $\mathcal{S}_{text}$;
return $\mathcal{P},\mathcal{N}$;
Algorithm 1 Automated Instance Annotation
Appendix D Annotation Data Quality Control
D.1 Validation Data Filtering
To establish a reliable upper bound on annotation quality, we construct a strategic audit subset rather than validating the full corpus. Algorithm 2 formalizes the sampling process as a greedy minimal-cover problem over observed US-GAAP concepts.
The algorithm takes as input the set of positive instances $\mathcal{P}$, together with the set of all taxonomy concepts observed in the corpus, $\mathcal{C}_{all}$. The goal is to produce a compact subset $\mathcal{P}^{*}$ that is both diverse and concept-complete.
Seeding for company diversity. We first iterate over companies and select up to $K$ instances per company. This prevents the audit set from being dominated by firms with longer filings or more frequent disclosures. After each selection, we update the set of uncovered concepts $\mathcal{C}_{rem}$ by removing concepts already represented among the seeds.
Greedy minimal cover. We then repeatedly add instances that contribute maximal marginal concept coverage. For each candidate instance, we compute the number of uncovered concepts it contains and select the instance with the largest score. When ties occur, we prefer shorter instances to reduce annotation effort while retaining conceptual value. After adding each instance, we update $\mathcal{C}_{rem}$ accordingly and continue until either all concepts are covered or no additional candidates remain.
Resulting audit subset. This procedure yields a compact subset that balances two competing objectives: coverage of the observed taxonomy and representation across companies. The resulting $\mathcal{P}^{*}$ therefore serves as a stress-test sample that is both manageable for human review and representative of the broader benchmark space.
Input: Positive instances $\mathcal{P}$ from Algorithm 1; Observed taxonomy concepts $\mathcal{C}_{all}$; Company cap $K=10$
Output: Strategic audit subset $\mathcal{P}^{*}$
Initialize: $\mathcal{P}^{*}\leftarrow\emptyset$;
Initialize: $\mathcal{C}_{rem}\leftarrow\mathcal{C}_{all}$ // Set of concepts yet to be covered;
Step 1: Diversity-Preserving Seeding
foreach unique company in $\mathcal{P}$ do
   Select up to $K$ instances belonging to this company and add to $\mathcal{P}^{*}$;
   Update $\mathcal{C}_{rem}$ by removing concepts covered by these seed instances;
Step 2: Greedy Minimal Cover
while $\mathcal{C}_{rem}$ is not empty and candidates remain in $\mathcal{P}$ do
 foreach candidate instance $i\in\mathcal{P}\setminus\mathcal{P}^{*}$ do
      Calculate $Score(i)=|\text{Concepts in }i\cap\mathcal{C}_{rem}|$;
  Select instance $i^{*}$ with the highest $Score(i)$;
 if multiple instances have the same $Score(i^{*})$ then
      Break ties by selecting the instance with the shortest text length;
  Add $i^{*}$ to $\mathcal{P}^{*}$ and update $\mathcal{C}_{rem}$ by removing newly covered concepts;
return $\mathcal{P}^{*}$;
Algorithm 2 Strategic Audit Set Sampling
D.2 Annotator Demography
To ensure reliable verification of financial entities and taxonomy mappings, all audits were conducted by annotators with prior exposure to XBRL-based reporting and financial statement analysis. The review team consisted of two junior analysts and one senior adjudicator.
The two junior analysts are graduate-level researchers with training in accounting and financial data analytics. Both have received formal instruction on US-GAAP concepts, XBRL tagging conventions, and common disclosure practices in SEC filings. They have prior experience working with financial datasets in academic or industry projects, including tasks involving concept normalization, document review, and regulatory text interpretation. Before participating in the audit, each analyst completed a guided calibration session using our Verification Guideline to ensure consistent treatment of numerical entities, metadata attributes, and taxonomy semantics.
The senior adjudicator has over a decade of experience in financial reporting analysis and structured disclosure systems. Their professional background includes work with XBRL instances, taxonomy structure, and validation procedures used in regulatory environments. In addition to overseeing the audit workflow, the senior adjudicator resolved disagreements, provided clarification on ambiguous tagging scenarios, and ensured alignment with US-GAAP definitions and XBRL best practices.
Together, this composition provides both annotation robustness and domain oversight: junior analysts contribute detailed review capacity, while the senior adjudicator ensures conceptual correctness and methodological consistency across the audit.
All annotators were members of the research team and were invited to participate as part of their normal research training and collaborative project responsibilities. Participation was voluntary, and no separate monetary compensation was provided. Because the work involved publicly available financial disclosures and did not require annotators to process personal or sensitive data, the review posed minimal risk to participants.
D.3 Validation Guideline
D.3.1 Task Definition
Given a context and an extracted triplet consisting of:
  • an entity,
  • its entity type,
  • a US-GAAP concept,
Your task is to determine whether the triplet is correct.
Make a binary decision:
  • 1 for correct,
  • 0 for incorrect.
D.3.2 Validation Procedure
Follow the rules below, the definitions of entity type are shown in Table 10:
  1. 1.
    Entity Check Determine whether the extracted entity is a numerical value.
    • If yes, proceed to Step 2.
    • If no, set is_correct = 0.
  2. 2.
    Entity Type Validation Verify whether the identified entity type is correct based on the context and the definitions in Table 10.
    • If yes, proceed to Step 3.
    • If no, set is_correct = 0.
  3. 3.
    US-GAAP Concept Validation Assess whether the assigned US-GAAP concept is appropriate based on the taxonomy tag definitions.
    • If yes, set is_correct = 1.
    • If no, set is_correct = 0.
Table 10: Entity type definitions used for validation.
Entity Type Definition
monetaryItemType Financial amounts expressed in currency, such as revenue, profit, or total assets.
integerItemType Counts of discrete items, such as the number of employees or total transactions.
perShareItemType Per-share values, such as earnings per share (EPS) or book value per share.
sharesItemType Counts of shares, such as outstanding shares or ownership stakes.
percentItemType Ratios or percentages, such as tax rates, growth rates, or discount rates, usually expressed with a percentage symbol (%).
D.3.3 Important Instructions
  1. 1.
    Non-Arabic Formats Financial numerical entities may appear in word form (e.g., ten million) and must be correctly identified and converted into standard numerical format.
  2. 2.
    Magnitude Terms If a number is followed by a magnitude term (e.g., hundred, million, billion), do not expand it into the full numerical value:
    • Two hundred $\rightarrow$ extract only two, not 200.
    • 10.6 million $\rightarrow$ extract only 10.6, not 10,600,000.
  3. 3.
    Standardization of Format Remove formatting symbols while preserving the numerical value:
    • Remove currency symbols (e.g., USD).
    • Remove percentage signs (e.g., %).
    • Remove commas (e.g., 1,000 $\rightarrow$ 1000).
D.4 Agreement Computation
We report both raw agreement and Cohen’s $\kappa$ to quantify inter-annotator reliability.
Let $N$ denote the total number of instances, and let $n_{yy}$ be the number of cases on which both annotators assign the same label $y$. The raw agreement is
$P_{o}=\frac{1}{N}\sum_{y}n_{yy}.$ (5)
To adjust for agreement occurring by chance, we compute Cohen’s $\kappa$. Let $p_{1y}$ and $p_{2y}$ denote the marginal probabilities that annotator 1 and annotator 2 assign label $y$, respectively. The expected chance agreement is
$P_{e}=\sum_{y}p_{1y}\,p_{2y},$ (6)
and Cohen’s $\kappa$ is
$\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}.$ (7)
Following landis1977measurement, our $\kappa=0.81$ indicates substantial agreement.
Appendix E Evaluation Metrics
To provide a fair evaluation of overall benchmark performance, we adopt a set of metrics, focusing primarily on macro-level and micro-level evaluation strategies inspired by the previous work sharma2023financial. Macro-level evaluation computes precision, recall, and F1 scores independently for each BIO-concept label derived from the US-GAAP taxonomy, and then averages them without weighting. This ensures that each concept, including rare or infrequent ones, contributes equally to the final score, making it especially suitable for domains with skewed label distributions. In contrast, micro-level evaluation aggregates token-level true positives, false positives, and false negatives across all labels before computing precision, recall, and F1. This approach emphasizes the model’s overall tagging accuracy by treating every token equally and thus better reflects performance on frequent concepts. Together, these two metrics provide a balanced view of both per-concept performance and overall tagging quality.
For the FinNI subtask, the objective is to extract correct (entity, type) pairs, that is $(\texttt{Fact},\texttt{Type})$, from the financial document. Let $\mathcal{G}={(e_{i},l_{i})}$ denote the set of ground-truth (entity, type) pairs, and $\mathcal{P}={(e^{\prime}_{i},l^{\prime}_{i})}$ denote the set of predicted (entity, type) pairs. We evaluate the performance based on the following metrics:
$\text{Precision}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{P}|}$ (8)
$\text{Recall}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{G}|}$ (9)
$\text{F1}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (10)
where $(e_{i},l_{i})=(e^{\prime}_{j},l^{\prime}_{j})$ if and only if both the entity span $e$ and its assigned type $l$ exactly match.
For the FinCL subtask, Given a set of queries $\mathcal{Q}=\{q_{1},q_{2},\dots,q_{N}\}$, where each $q_{i}$ is associated with a ground-truth concept $c_{i}^{*}$ and a predicted concept $\hat{c}_{i}$, the accuracy is defined as:
$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\hat{c}_{i},c_{i}^{*})$ (11)
where $\delta(\hat{c}_{i},c_{i}^{*})=1$ if $\hat{c}_{i}=c_{i}^{*}$, and 0 otherwise.
Appendix F Evaluation Models Details
Table 11 provides an overview of the models evaluated in this study, categorized by openness, domain specialization, and architectural foundation. The evaluation covers a diverse range of models:
  • Closed-source LLMs: We include GPT-4o hurst2024gpt, accessed via OpenAI’s API, as a representative of cutting-edge proprietary models with demonstrated performance across a variety of NLP tasks. Although model size details are undisclosed, GPT-4o serves as an upper-bound reference in our benchmark.
  • Open-source LLMs: This group encompasses recent, high-performing open models such as DeepSeek-V3 (685B) liu2024deepseek, DeepSeek-R1-Distill-Qwen (32B) deepseekai2025deepseekr1incentivizingreasoningcapability, and multiple variants of Qwen3  qwen3technicalreport(ranging from 0.6B to 32B). We also include Llama-3.3, Llama-3.2, and 3.1 variants  grattafiori2024llama (70B, 3B, and 8B), Llama-4-Scout-17B-16E-Instruct meta2025llama(109B), as well as Google’s gemma-2-27B  gemma_2024, to ensure architectural diversity and scalability comparison. These models are primarily instruction-tuned and optimized for general-purpose NLP tasks.
  • Financial-specific LLMs: We evaluate Fino1-8B  qian2025fino1, a domain-specialized model trained on financial corpora, designed to better capture the terminology and structure unique to financial disclosures. This category allows us to assess the benefits of domain adaptation in complex tagging and reasoning tasks.
  • Pretrained Language Models (PLMs): To establish strong baselines, we include non-generative encoder models: BERT-large  DBLP:journals/corr/abs-1810-04805, FinBERT  araci2019finbert, and SECBERT  loukas2022finer. These models have been widely used in prior financial NLP tasks and allow for a comparative analysis between generative LLMs and traditional pretrained models in terms of domain understanding and structured output capability.
Together, these models offer a comprehensive evaluation spectrum, from general-purpose to domain-specific, encoder-based to decoder-based, and open to closed source, facilitating an in-depth assessment of their performance across our proposed benchmark tasks.
Table 11: Model categories and corresponding repositories.
Model Size Source
Closed-source Large Language Models
GPT-4o - gpt-4o-2024-08-06
Open-source Large Language Models
DeepSeek-V3 685B deepseek-ai/DeepSeek-V3
Llama-4-Scout-17B-16E-Instruct 109B meta-llama/Llama-4-Scout-17B-16E-Instruct
Llama-3.3-70B-Instruct 70B meta-llama/Llama-3.3-70B-Instruct
DeepSeek-R1-Distill-Qwen 32B deepseek-ai/DeepSeek-R1-Distill-Qwen-32B
Qwen3-32B 32B Qwen/Qwen3-32B
gemma-2-27b-it 27B google/gemma-2-27b-it
Qwen3-14B 14B Qwen/Qwen3-14B
Llama-3.1-8B-Instruct 8B meta-llama/Llama-3.1-8B-Instruct
Llama-3.2-3B-Instruct 3B meta-llama/Llama-3.2-3B-Instruct
Qwen3-1.7B 1.7B Qwen/Qwen3-1.7B
Qwen3-0.6B 0.6B Qwen/Qwen3-0.6B
Financial-specific Large Language Models
Fino1 8B TheFinAI/Fino1-8B
Pretrained Language Models
BERT-large $\sim$340M google-bert/bert-large-uncased
FinBERT $\sim$110M ProsusAI/finbert
SECBERT $\sim$110M nlpaueb/sec-bert-base
Appendix G The details for the fine-tuning PTMs
G.1 Training data collection and processing
Similar to the collection process for the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark data, we gathered an additional 10 annual 10-K financial reports filed with the SEC for the period from February 13, 2024, to February 13, 2025, as summarized in Table 12. These reports contain a total of 33,848 standard taxonomy-tagged facts. Using BeautifulSoup to parse these documents, we identified 22,847 narrative sentences (approximately 5.5 million characters) and 1,236 financial tables. The companies included in this dataset follow the XBRL standard, ensuring comprehensive coverage for training PTMs.
Table 12: Financial report statistics summary for raw training data
Item Information
Report type 10-K
Period 2024-02-13 to 2025-02-13
#Company 10
#Sentence 22,847
#Table 1,236
#Characters 5,539,198
#Standard Tags 33,848
After collection, we employed the same procedure to filter texts and tables, subsequently annotating numerical entities, entity types, and US-GAAP tags (concepts). Finally, as detailed in Table 13, we generated a total of 1,116 sentences and 953 tables as the training set for PTMs. Specifically, the sentence-level data consists of 558 positive and 558 negative instances, averaging approximately 84.24 tokens ($\pm$ 69.29), with 1.22 annotated entities and concepts per sentence. The table-level data comprises 594 positive and 359 negative instances, with a significantly higher average of 1,281.86 tokens ($\pm$ 6,438.37), and approximately 25 entities and concepts annotated per table. Overall, the annotated dataset includes 25,199 entities, covering 1,435 unique US-GAAP concepts.
Table 13: Statistics of training data (tokens calculated with “cl100k_base” tokenizer, $\pm$ standard deviation).
Structure Pos/Neg #Instance Avg. Tokens/S Avg. Entities/S Avg. Concepts/S Total Entities Unique Concepts
Sentence Positive 558 $84.24\pm 69.29$ $2\pm 1.78 1.22±1$$2\pm 1.78$ 25,199 1,435
Negative 558
Table Positive 594 $1281.86\pm 6438.37$ $25.00\pm 213.77$ $25.00\pm 213.77$
Negative 359
However, to align with the extreme classification format used in previous XBRL tagging benchmarks, we directly adopt the US-GAAP tags as entity labels, annotating each token in sentences and tables using the BIO scheme. Specifically, B denotes the beginning of an entity phrase, I marks the continuation (inside) of an entity phrase, and O indicates tokens outside of any entity. As shown in Figure 8, “4.9” and “4.5” are single-token numerical entities labeled only with a B prefix (e.g., “B-us-gaap:AccountsReceivableNetNoncurrent”). To comprehensively cover all US-GAAP tags, we combine the entire set of 17,688 tags from the US-GAAP 2023 and 2024 taxonomy with the BIO labeling scheme to construct an extreme classification label space, resulting in 34,777 unique entity labels ($2\times 17388+1$).
Figure 8: An example for training set annotation with BIO scheme.
After constructing the training set, we reconstruct the testing set from the original benchmark dataset. The training settings are detailed below.
G.2 Training settings
We fine-tune three pretrained models, BERT-large DBLP:journals/corr/abs-1810-04805, FinBERT araci2019finbert, and SECBERT loukas2022finer, on our training set using the HuggingFace Transformers library. All models are trained with a batch size of 4, a learning rate of 3e-5, and for 20 epochs. Optimization is performed using AdamW without gradient accumulation or early stopping. Token classification heads are randomly initialized and trained jointly with the base encoder. Input sequences are tokenized with a maximum length of 512, and labels are aligned at the sub-token level following the BIO tagging scheme. Loss is computed only on the first sub-token of each word to avoid misalignment bias.
Training is conducted on two NVIDIA A5000 GPUs (24GB each) using data parallelism for 24 hours. All other hyperparameters follow the default settings in the HuggingFace Trainer API. Models are evaluated using the checkpoint from the final epoch. All experiments are run under a fixed random seed to ensure reproducibility.
Appendix H Retrieval Results in FinCL subtask
We investigate the impact of different context construction strategies for the queried entity at the retrieval stage. We consider two approaches: Fixed-Window Context (FWC) and Structure-Aware Context (SAC). In the FWC strategy, context is constructed by extracting a fixed window of 50 characters before and after the entity mention, regardless of whether the entity appears in a sentence or a table. In contrast, the SAC strategy builds context based on the structural location of the entity: if the entity appears in a sentence, the entire sentence is used; if the entity appears in a table, we linearize the entire row into a Markdown-style key-value format to serve as the context.
Table 14: Acc@k retrieval performance on the FinCL task. FWC: fixed window-based; SAC: structure-aware.
Strategy Structure Acc@1 Acc@10 Acc@20 Acc@50 Acc@100 Acc@150 Acc@200
FWC Sentence 0.0658 0.2614 0.3562 0.4650 0.5618 0.6114 0.6502
Table 0.0000 0.0029 0.0038 0.0042 0.0050 0.0065 0.0089
Overall 0.0055 0.0237 0.0316 0.0409 0.0492 0.0544 0.0608
SAC Sentence 0.0696 0.2742 0.3631 0.4798 0.5701 0.6125 0.6535
Table 0.0159 0.0872 0.1245 0.1938 0.2452 0.2723 0.2959
Overall 0.0202 0.1152 0.1534 0.2188 0.2727 0.3012 0.3274
From Table 14, we observe that the SAC strategy consistently outperforms FWC, particularly in the table context, highlighting the importance of aligning the context window with the underlying structural unit.
For sentence-based entities, both strategies yield comparable results, with SAC achieving slightly higher Acc@1 (0.0696 vs. 0.0658) and showing marginal improvements across all cutoff values (e.g., Acc@100 of 0.5701 vs. 0.5618). This suggests that even for relatively unstructured text, preserving sentence boundaries provides minor benefits over fixed-length windows.
In contrast, for table-based entities, the performance gap is substantial. SAC achieves significantly higher retrieval accuracy (e.g., Acc@100 of 0.2452 vs. 0.0050), indicating that row-level context is far more informative than arbitrary character windows when dealing with tabular structures. FWC performs poorly in this setting, likely due to the fragmented and semantically sparse nature of partial table text.
When aggregating results across both structures, SAC outperforms FWC by a wide margin at all retrieval depths (e.g., Acc@200 of 0.3274 vs. 0.0608). These findings underscore the importance of structure-aware context construction, especially in scenarios where inputs span multiple formats such as sentences and tables.
Appendix I Error cases
To better illustrate the semantic ambiguity challenge, we present a representative error case from our best-performing model, DeepSeek-V3. Consider the following input paragraph:
“Cash equivalents include term deposits with banks, money market funds, and all highly liquid investments with original maturities of three months or less…… At December 28, 2024, we had restricted cash of \$31 million recorded in other current assets and restricted cash of \$121 million recorded in other non-current assets.”
This paragraph contains 2 “monetary entities”: 31 and 121. The gold concepts and predicted concepts that are assigned to both entities are shown in Table 15. For case 1, the model’s choice was likely influenced by the phrase “recorded in current assets”. Yet the paragraph describes the overall composition of restricted cash, and the correct concept reflects this aggregate view rather than a single classification. For case 2, although both concepts denote non-current restricted cash, the model selected a narrower variant, showing the difficulty of distinguishing between highly similar concepts with subtle differences in scope.
Table 15: Error cases from DeepSeek-V3 rerank.
Case 1: \$31m
Gold Concept us-gaap:RestrictedCashAndCashEquivalentsAtCarryingValue
Gold Concept Rank 18 / 200 (not in Top-5)
Model Prediction us-gaap:RestrictedCashAndCashEquivalentsCurrent
Case 2: \$121m
Gold Concept us-gaap:RestrictedCashAndCashEquivalentsNoncurrent
Gold Concept Rank 2 / 200 (in Top-5)
Model Prediction us-gaap:CashCashEquivalentsRestrictedCashAndRestrictedCashEquivalents
Prompt Template for FinNI Subtask You are a financial information extraction expert specializing in identifying financial numerical entities in XBRL reports. Your task is to extract all such numerical entities from the provided text or serialized <table></table> data and classify them into one of five categories: - "integerItemType": Counts of discrete items, such as the number of employees or total transactions. - "monetaryItemType": Financial amounts expressed in currency, such as revenue, profit, or total assets. - "perShareItemType": Per-share values, such as earnings per share (EPS) or book value per share. - "sharesItemType": Counts of shares, such as outstanding shares or ownership stakes. - "percentItemType": Ratios or percentages, such as tax rates, growth rates, or discount rates, usually expressed with a percentage symbol ("%"). Important Instructions: (1) Financial numerical entities are not limited to Arabic numerals (e.g., 10,000). They may also appear in word form (e.g., "ten million"), which must be correctly identified and converted into standard numerical format. (2) Not all numbers in the text should be extracted. Only those that belong to one of the five financial entity categories above should be included. Irrelevant numbers (such as phone numbers, dates, or general IDs) must be ignored. (3) If a number is followed by a magnitude term (e.g., Hundred, Thousand, Million, Billion), do not expand it into the full numerical value. * "Two hundred" -> Extract only "two", not "200". * "10.6 million" -> Extract only "10.6", not "10,600,000". (4) Standardize numerical formatting by removing currency symbols (e.g., "USD"), percentage signs ("%"), and commas (",") while preserving the numeric value. These elements must be removed to ensure consistency. (5) Output the extracted financial entities in JSON list format without explanations, structured as follows: {"result":[{"Fact": <Extracted Numerical Entity>, "Type": <Identified Entity Type>}]} Input: {text/table} Output:
Figure 9: Prompt template used for the FinNI subtask.
Prompt Template for FinCL Subtask (Reranking) You are a financial tagging assistant trained in US-GAAP taxonomy. Given a query consisting of an entity, its type, its surrounding context, and the source format (either text or table), your task is to select the single most appropriate US-GAAP tag from a list of 200 candidate tags. Make your decision by carefully analyzing the meaning and context of the entity and matching it with the semantics of the tags. Only output one tag, the best match. Do not explain or list multiple tags. The output is a JSON format: {"result": <the best matched tag>}. Input Query: <entity> + <entity type> + <context> Candidate Tags: {Top 200 US-GAAP tags} Answer:
Figure 10: Prompt template used for the Reranking stage in the FinCL subtask.
Prompt Template for Ablation You are an XBRL tagging expert specializing in annotating financial numerical facts in XBRL reports. Your task is to (1) extract all such numerical entities from the provided text or serialized <table></table> data, (2) classify them into one of five categories, and (3) assign an appropriate US-GAAP tag to each entity. Categories: - "integerItemType": Counts of discrete items, such as the number of employees or total transactions. - "monetaryItemType": Financial amounts expressed in currency, such as revenue, profit, or total assets. - "perShareItemType": Per-share values, such as earnings per share (EPS) or book value per share. - "sharesItemType": Counts of shares, such as outstanding shares or ownership stakes. - "percentItemType": Ratios or percentages, such as tax rates, growth rates, or discount rates, usually expressed with a percentage symbol ("%"). US-GAAP tags: - A US-GAAP tag is a standardized semantic label used in XBRL filings to identify specific financial concepts defined by the U.S. Generally Accepted Accounting Principles (GAAP). Each tag represents a distinct accounting item and enables consistent, machine-readable financial reporting. - Examples: "us-gaap:AssetsCurrentAbstract", "us-gaap:AccruedInsuranceNoncurrent". Important Instructions: (1) Financial numerical entities are not limited to Arabic numerals (e.g., 10,000). They may also appear in word form (e.g., "ten million"), which must be correctly identified and converted into standard numerical format. (2) Not all numbers in the text should be extracted. Only those that belong to one of the five financial entity categories above should be included. Irrelevant numbers (such as phone numbers, dates, or general IDs) must be ignored. (3) If a number is followed by a magnitude term (e.g., Hundred, Thousand, Million, Billion), do not expand it into the full numerical value. * "Two hundred" -> Extract only "two", not "200". * "10.6 million" -> Extract only "10.6", not "10,600,000". (4) Standardize numerical formatting by removing currency symbols (e.g., "USD"), percentage signs ("%"), and commas (",") while preserving the numeric value. These elements must be removed to ensure consistency. (5) You should assign the most appropriate US-GAAP tag to each identified entity based on your internal understanding of the 2024 US-GAAP taxonomy. (6) Output the extracted financial entities in JSON list format without explanations, structured as follows: {"result":[{"Fact": <Extracted Numerical Entity>, "Type": <Identified Entity Type>, "Tag": <Assigned US-GAAP tag>}]} Input: {text/table} Output:
Figure 11: Prompt template used for ablation study.